Could you please help me with this problem? I came up with 42/91 but I don't know if that is right because I don't know if I did them in the right order.
2 1/3 / 1 1/6 * 3 1/4 / 1 1/3

Answer by stanbon(75887)

I came up with 42/91 but I don't know if that is right because I don't know if I did them in the right order.
2 1/3 / 1 1/6 * 3 1/4 / 1 1/3
---
[(7/3)/(7/6)]
----
= (6/3)*(39/16)
----
= (1/2)(39/16)
---
= (39/32)
============
Cheers,
Stan H.
===============